OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16)

OpenStudy (p0sitr0n):

same slope, but different b. 16=2+b 14=b y = x + 14

OpenStudy (anonymous):

These are the choices A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

b and d are the same thing.......

OpenStudy (anonymous):

so is it one of them?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@RoseDryer can you help me

OpenStudy (raden):

looks there are no options be right

OpenStudy (anonymous):

@RadEn is right. @P0sitr0n is correct when they said y= x+14 I would suggest telling your teacher that there are no incorrect answers...

OpenStudy (anonymous):

What should I do?

OpenStudy (anonymous):

Well if you can email a teacher do that if not just guess because technically no matter which one you choose it'll be wrong.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is it one of them?